OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle? 1 23 square 265 square 23

OpenStudy (anonymous):

\[\sqrt{23}\]

OpenStudy (anonymous):

can you tell me why

OpenStudy (anonymous):

do you know Pythagoras theorem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so H^2= B^2+P^2 12^2= 11^2+P^2 P^2=12^2-11^2 undrtP=undert23 you get it now ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

great :D

OpenStudy (anonymous):

let x be the length of the third side. \[12^{2} = 11^{2} + x ^{2}\] \[x ^{2} = 12^{2} - 11^{2}\] \[x = \sqrt{144 - 121}\] \[x = \sqrt{23}\]